40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Wed Mar 23 01:14:27 2011, Zach, Laser, GYRO, signal routing finished, gyro locked table.pngscope.pngrack.pnglid_hole.png
    Reply  Thu Mar 24 09:07:17 2011, Alastair & Zach, Laser, GYRO, RFAM re-measured full_spectrum_RFAM.png5_MHz_to_35_MHz.png
       Reply  Thu Mar 24 11:35:38 2011, rana, Laser, GYRO, RFAM re-measured 
Message ID: 1359     Entry time: Wed Mar 23 01:14:27 2011     Reply to this: 1362
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: signal routing finished, gyro locked 

 [Alastair, Zach]

We finished re-setting up the gyro this evening. Alastair cut a hole in the top of the side of the IO box so that the lid can be put on with the necessary cables fed in. Everything, including the PLL, is back online with everything working from the rack.

I had some of the time-varying error signal offset issue coming back, but I was able to get rid of it by iteratively adjusting the HWPs before and after the EOM so that 1) there was minimal RFAM and 2) the power was balanced between the CCW/CW beams. I also adjusted the modulation frequency again slightly because it looked as if we were sitting on a bad spot with respect to the modulator's AM/PM response. We still need to take a proper transfer function of this (with better resolution than the old one) to be sure, and once we're done settling on a frequency we have to retune the REFL PDs and EOM resonant circuit.

I will continue with the loop characterization tomorrow.

table.pngscope.pngrack.pnglid_hole.png

ELOG V3.1.3-